Please help me with the following problem:
If we extablish an identity:
(cos^2 A)/(1 + sin A) would equal which of the following?
1. cos^2 A + cot A cos A
2. 1 + sin A
3. 1 - sin A
4. 1
Thank you, Lennie
: Please help me with the following problem:
If we extablish an identity:
(cos^2 A)/(1 + sin A) would equal which of the following?
1. cos^2 A + cot A cos A
2. 1 + sin A
3. 1 - sin A
4. 1
Thank you, Lennie

Answer by jim_thompson5910(9217)

Start with the given expression.


Rewrite as


Factor to get


Photobucket - Video and Image HostingCancel out the common terms.

Simplify


So simplifies to .


In other words, where

Question 150175: Please help me with the following problem:
If we extablish an identity:
(cos^2 A)/(1 + sin A) would equal which of the following?
1. cos^2 A + cot A cos A
2. 1 + sin A
3. 1 - sin A
4. 1
Thank you, Lennie
: Please help me with the following problem:
If we extablish an identity:
(cos^2 A)/(1 + sin A) would equal which of the following?
1. cos^2 A + cot A cos A
2. 1 + sin A
3. 1 - sin A
4. 1
Thank you, Lennie

Answer by stanbon(18788)

If we extablish an identity:
(cos^2 A)/(1 + sin A) would equal which of the following?
-----------------
cos^2 A = 1 - sin^2 A
-----------------
Your Problem:
cos^2 A / (1 + sin A)
= (1 - sin^2 A)/(1+sin A)
= [(1-sin A)(1+sin A)]/(1+sin A)
= 1 - sin A
-------------------------
Cheers,
Stan H.

1. cos^2 A + cot A cos A
2. 1 + sin A
3. 1 - sin A
4. 1